OpenStudy (anonymous):

14. Sharon earns $25 per item she sells plus a base salary of $100 per week. Write and solve an inequality to find how many items she must sell to earn at least $700 per week. Provide your conclusion as a complete sentence.

OpenStudy (anonymous):

25x+100=700 25x =700-100 x =(700-100)/25=24 so Sharon needs to sell 24 items each week to earn 700$

OpenStudy (anonymous):

thank you